OpenStudy (anonymous):

9x-(7x+1/2)=3x+(1/2-X)

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

do you know how to do the distributive property?

OpenStudy (anonymous):

Yea bt im nt sure hw to strt

OpenStudy (anonymous):

9x-(7x+1/2)=3x+(1/2-X) 9x-7x=2x 2x+1/2=3x+1/2-x 2x+1/2=2x+1/2 2x=2x x=x

OpenStudy (anonymous):

the answer is "all real numbers"

OpenStudy (anonymous):

\[9x -1(7x+\frac{ 1 }{ 2 }) = 3x + 1(\frac{ 1 }{ 2 }-x)\] does that look nicer? :}

OpenStudy (anonymous):

the answer is x=x not just all real numbers

OpenStudy (anonymous):

Thanks can u help with thid

OpenStudy (anonymous):

-4/7x=-12+4

OpenStudy (anonymous):

well to get rid of the fraction multiply both sides by 7 to get what

OpenStudy (anonymous):

7x sorry

OpenStudy (anonymous):

or divide each side by four leaving the left side negative seven

OpenStudy (anonymous):

7 cancels out?

OpenStudy (anonymous):

\[\frac{ -4 }{ 7x } = -12 + 4\]

OpenStudy (anonymous):

9x-(7x+1/2)=3x+(1/2-X) lets say x=34 306-238+1/2=102+1/2-34 68+1/2=68+1/2 - lets say that x=113 1017-791+1/2)=339+1/2-113 226+1/2=226+1/2 i can do this all day, because X= ALL REAL NUMBERS

OpenStudy (anonymous):

on the left do you want to get rid of the seven or the four

OpenStudy (anonymous):

7

OpenStudy (anonymous):

i got x = 14

OpenStudy (anonymous):

Hw/

OpenStudy (anonymous):

multiply both sides by 7x to get \[-4 = -84x + 28x\] \[-4 = -56x\] \[0 = \frac{ -56x }{ -4 }\] 0 = 14x 0 = x

OpenStudy (anonymous):

@lharrell97 if x=x then it can be anything because you don't know the value of x so it could be an imaginary number or it could be elephant...not JUST all real numbers

OpenStudy (anonymous):

Thanks bt i hve more x+5 over 3 = x-3 over 4

OpenStudy (anonymous):

because you dont know the value of X, it could be any possible number. x could equal 3 billion or it could equal 2. the answer therefore is X= all real numbers... although i can see where you're coming from haha

OpenStudy (anonymous):

Hw would i work it out?

OpenStudy (anonymous):

i think it's probable that X=elephant probably...yeah probably lol

OpenStudy (anonymous):

look like that? \[\frac{ x + 5 }{ 3 } =\frac{ x - 3 }{ 4 }\]